OpenStudy (anonymous):

arccos(cos(94pi/59))

OpenStudy (anonymous):

HINT: \[\cos^{-1} (\cos x) = x\]

OpenStudy (anonymous):

arccos(cos(94pi/59)) = 94pi/59

zepdrix (zepdrix):



OpenStudy (anonymous):

94pi/59 is in the 4th quad. so... 94pi/59 -pi = 35/59 pi

OpenStudy (anonymous):

We know the range of the arc cosine function is limited to [0,pi], right? We know that 90pi/59 is pretty close to 90pi/60 or 3pi/2. The cos(3pi/2) is close to 0 The arccos(0) is pi/2 The reference angle for the 90pi/59 would be in quad 4 and measure 28/59 pi. SO I'm thiniking the answer to the problem would be 28pi/59